OpenStudy (anonymous):

Can someone help me please!? Evaluate the piecewise-defined function for the given values. If f(x) = x for x ≥ 0 x2 for x < 0 , find f(5), f(6), f(−2), and f(−3).

OpenStudy (anonymous):

x2 = x^2

OpenStudy (anonymous):

@ sm101 for x>=0 what is f(x)=? then what is f(5)=? and f(6)=?

OpenStudy (anonymous):

I don't know... ??

OpenStudy (anonymous):

for x>=0 what is f(x)=? it is f(x)=x just use the formula given then what is f(5)=? and f(6)=?.... f(5)=x and f(6)=6..

OpenStudy (anonymous):

next f(x)=x^2 for x<o use this formula for x<0 w/c is x=-2 and x=-3 therefore f(−2)=___?), and f(−3).=___?

OpenStudy (anonymous):

ok.. I'll try

OpenStudy (anonymous):

ok good just plug into the formula f(x)=x^2 right?

OpenStudy (anonymous):